A father is 27 years older than his son..In 3 years he will be twice as old as his son. Calculate the age of the father and the son.
Answer by macston(5194)

F=father's age; S=son's age
F=S+27
F+3=2(S+3) Substitute for F
S+27+3=2S+6
S+30=2S+6 Subtract S from each side.
30=S+6 Subtract 6 from each side.
24=S ANSWER 1: The son is 24 years old
F=S+27=24+27=51 ANSWER 2: The father is 51 years old.
CHECK IN three years, the father will be twice as old as the son:
F+3=2(S+3)
51+3=2(24+3)
54=2(27)
54=54